40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 8 23:59:42 2013, yuta, Update, Locking, PR2-flipped half-PRC mode scan halfPRCmodescan.png
    Reply  Sat Feb 9 11:25:35 2013, Koji, Update, Locking, PR2-flipped half-PRC mode scan 
       Reply  Sun Feb 10 17:30:39 2013, yuta, Update, Locking, PR2-flipped half-PRC mode scan modescan_pitmisalign.pngmodescan_yawmisalign.png
          Reply  Mon Feb 11 13:15:16 2013, yuta, Update, Locking, PR2-flipped half-PRC mode scan 
          Reply  Mon Feb 11 21:03:15 2013, yuta, Update, Locking, PR2-flipped half-PRC mode scan modespacing_pit.pngmodespacing_yaw.png
Message ID: 8049     Entry time: Fri Feb 8 23:59:42 2013     Reply to this: 8050
Author: yuta 
Type: Update 
Category: Locking 
Subject: PR2-flipped half-PRC mode scan 

I did mode scan of PR2-flipped half-PRC to see if it behaves as we expect.
Measured finesse was 107 +/- 5 and g-factor is 0.98997 +/- 0.00006.
g-factor is 0.9800 +/- 0.0001.  (Edited by YM; see elog #8056)

 
Finesse tells you that we didn't get large loss from flipped PR2.
Since we have convex TM in front of BS, PRC will be more stable than this half-PRC.

Method:
 1. Aligned half-PRC using input TT1 and TT2 by maximizing POP DC during lock. It was not so easy because POP DC fluctuates much at ~ 3 Hz with amplitude of ~ 30 % of the maximum value because of the beam motion (movie on  elog #8039).

 2. Unlocked half-PRC and took POP DC and PRC error signal data;

> /opt/rtcds/caltech/c1/scripts/general/getdata -d 1 -o /users/yuta/scripts/PRCmodescan C1:LSC-POPDC_OUT C1:LSC-REFL11_I_ERR

  Ran again and again until I get sufficiently linear swing through upper/lower sidebands.

 3. Ran modescan analyzing scripts (elog #8012).

Result:
 Below is the plot of POP DC and PRCL error signal (REFL11_I).
halfPRCmodescan.png

 By averaging 5 sets of peaks around TEM00;

Time between TEM00 and sideband  0.0347989  pm  0.00292257322372  sec
Calibration factor is  317.995971137  pm  26.7067783894  MHz/sec
FSR is  34.5383016129  MHz
FWHM is  0.323979022488  pm  0.0145486106353  MHz
TMS is  1.55827297374  pm  0.00439737672808  MHz
Finesse is  106.606598624  pm  4.78727876459
Cavity g-factor is  0.989971692098  pm  5.65040851566e-05
Cavity g-factor is  0.980043951156  pm  0.000111874889586

Discussion:
 Measured finesse is similar to measured PRM-PR2 cavity finesse(108 +/- 3, see elog #8012). This means loss from flipped PR2 and beam path from PR2 to TM is small.

 I'm a little suspicious about measured g-factor because it is hard to tell which peak is which from the mode scan data. Since half-PRC was not aligned well, high HOMs may contribute to POP DC. Astigmatism also splits HOM peaks.

 PRC 3 Hz beam motion was there for long time (see, for example, elog #6954). BS is unlikely to be the cause because we see this motion in half-PRC, too.
 Also, beam spot motion was not obvious in the PRM-PR2 cavity. My hypothesis is; stack resonance at 3 Hz makes PR2/PR3 angular motion and folding by PR2/PR3 makes the beam spot motion.

Next things to do:
 * PRC g-factor
   - Calculate expected half-PRC g-factor with real measured curvatures, with error bar obtained from RoC error and length error (JAMIE)
   - Calculate expected PRC g-factor using measured half-PRC g-factor (JAMIE)
 * PRC 3 Hz beam motion
   - Do we have space to put oplevs for PR2/PR3?
   - Can we fix PR2/PR3 temporarily?
 * PRMI
   - Align incident beam, BS, REFL, AS, and MI using arms as reference
   - lock PRMI
   - PRC mode scan

ELOG V3.1.3-